### v1

@@ 3 Inspecting Pipelines
PyTerrier transformers can be defined with as little as a function that maps one relation to another. This flexility makes implementing transformers simple and allows arbitrary pipelines to be constructed, but can result in incompatible pipelines. Without knowing the input and output specifications of individual transformers, it can be challenging to know whether a pipeline is compatible. Previously the only way to check these would be to check the documentation or run an example through and check the outputs (or interpret the resulting error).
@@ 4 Visualizing Pipelines with Schematics
As shown in Figure 1(a), pipeline definitions can get complex, especially when they involve multiple components and stages. Therefore, we added a new pipeline visualization feature that renders pipelines as interactive HTML snippets. We call these visitations schematics. Schematics render the pipeline as a sequence of transformers and the data between then, allowing users to better understand the transformers in a pipeline and how data flows between them. Hovering over elements provide additional information, like specific information about columns or transformer settings. Schematics leverage the inspection features covered in Section 3. A non-interactive screenshot of a schematic is shown in Figure 1(b).

### v2

@@ 3 Inspecting Pipelines
PyTerrier transformers can be defined with as little as a function that maps one relation to another. This flexibility makes implementing transformers simple and allows arbitrary pipelines to be constructed, but can result in incompatible pipelines. Without knowing the input and output specifications of individual transformers, it can be challenging to know whether a pipeline is compatible. Previously the only way to check these would be to check the documentation or run an example through and check the outputs (or interpret the resulting error).
@@ 4 Visualizing Pipelines with Schematics
As shown in Figure 1(a), pipeline definitions can get complex, especially when they involve multiple components and stages. Therefore, we added a new pipeline visualization feature that renders pipelines as interactive HTML snippets. We call these visitations schematics. Schematics render the pipeline as a sequence of transformers and the data between then, allowing users to better understand the transformers in a pipeline and how data flows between them. Hovering over elements provides additional information, like specific information about columns or transformer settings. Schematics leverage the inspection features covered in Section 3. A non-interactive screenshot of a schematic is shown in Figure 1(b).